OpenStudy (hopeless):

There 429 people at a play. Admission was $8 each for adults and $4.50 each for childern. The total receipts were $2641. How many adults and how many children attended?

OpenStudy (anonymous):

226 children Solve the following for a and c: \[\left\{a 8+c \frac{45}{10}=2641,a+c=429\right\}\]\[\{a=203,c=226\} \]